OpenStudy (anonymous):

Each car in a rental fleet is either red or white, and there are 75 cars total. There are three more white than twice the number of red cars. How many cars are white? A) 24 B) 47 C) 48 D) 51

OpenStudy (anonymous):

\[75=2x+3 ?\]

OpenStudy (anonymous):

But I'm trying to find the value of white, my sister thinks its r2+3=w

OpenStudy (anonymous):

yes your sister is right. the only problem is you cannot solve an equation for two variables.

OpenStudy (anonymous):

the equations still not working :|

OpenStudy (anonymous):

the equation would be 75=r+w and 2r+3=w you will use substitution to solve for r and then you can solve for w.

OpenStudy (anonymous):

how do I find something in place of r?

OpenStudy (anonymous):

you can change the second equation to (w-3)/2=r and then you can substitute that for r to solve for w.

OpenStudy (anonymous):

you will get a new equation of 75=(w-3)/2+w solve for w.

OpenStudy (anonymous):

I think it is easier if you solve for r and then solve for w. because you don't have to work with parentheses.